Martin needs to mix some 40% alcohol solution with some 90% alcohol solution to obtain 15 liters of an 80% solution. How much of the 40% solution and how much of the 90% solution should he use?
I need the answer quickly, so it would be greatly appreciated if I could receive an answer within a few hours. Thank you very much.

Answer by mananth(16946)

percent ----------------quantity
Alcohol 40 ----------------x
Alcohol 90 ----------------15-x
Total 80.00% ----------------15
40x+90(15-x)=80 *15

40x+1350-90x=1200
40x-90x =1200-1350
-50x=-150
/-50
x=3 liters 40.00%
12 liters 90.00%